Bottles R Us Bottling Company provides the following information about its single product.

Targeted operating income $654,870
Selling price per unit $7.50
Variable cost per unit $4.60
Total fixed cost $163,560

How many units must be sold to earn the targeted operating income?

67,639

56,400

282,218

225,817
Answer #1

Contribution margin=Sales-Variable cost

=7.50-4.60=$2.9 per unit

Target Contribution margin=Fixed expenses+Target income

=163,560+654,870=$818430

Hence units to be sold=818430/2.9

=282,218 units(Approx).